# Better Gaussian Mechanism using Correlated Noise

### v1

@@ 4.1 Bounded density
Here we consider a setting where each data point contributes to a bounded number of queries. This setting was recently studied by [JY24]. They gave efficient constructions for adding elliptical Gaussian noise for two classes of linear queries. The relevant query for this paper is the Count problem, which is equivalent to the setting from Section 3 with the additional constraint that each data point has at most $k\leq d$ non-zero entries. Similar to our mechanism they add additional noise along ${\bf 1}^{d}$. They give closed-form solutions for the shape of the optimal ellipse minimizing expected squared $\ell_{2}$ norm. However, their result only applies when $k\leq d/2$. The setting we considered in Section 3 corresponds to the case where $k=d$. A natural question is whether the error can be improved for the range $d/2<k<d$. It turns out that there is almost no room for improvement. In fact, our mechanism is almost optimal for any $k\geq d/2$. The difference between the variance of our mechanism and the optimal ellipse of [JY24] with parameter $k=\lfloor d/2\rfloor$ is less than $1$ for any $d$.
In Figure 3 we recreated a plot from Figure 1 of [JY24]. They plot the error relative to the standard Gaussian mechanism. The improvement increases with $k$, but their mechanism is undefined for $k>d/2$. We extended the line with our mechanism for $k>d/2$ showing the improvement for all values of $k$. We also all three mechanisms without normalizing the variance. From the plot it is clear that their mechanism should be used for small $k$ while our mechanism is preferred for any $k>d/2$. As a small note we point out that their mechanism does not include an estimate of the dataset size. The optimal ellipse might differ slightly if we want to return an estimate of $n$ similar to Algorithm 1.
@@
Figure 3: Comparison of error for the bounded Count problem for $d=1000$. Note that the optimal Count ellipse by [JY24] requires $k\leq d/2$. The left plot corresponds to the right plot of [JY24, Figure 1] where we included our mechanism for $k>d/2$.
@@ 5 Related work
The Gaussian mechanism satisfies several definitions of differential privacy. Tight bounds are known for $(\varepsilon,\delta)$-DP [BW18], $(\alpha,\varepsilon)$-RDP [Mir17], $\rho$-zCDP [BS16], and $\mu$-GDP [DRS22]. Our privacy guarantees follow from a reduction to the standard Gaussian mechanism. As such, our results apply to any of the above definitions of differential privacy. We choose to present our mechanism in terms of $\mu$-GDP because this definition exactly describes the privacy guarantees of the Gaussian mechanism. For all problems we consider in this paper our privacy guarantees for worst-case pairs of neighboring datasets exactly match those of the standard Gaussian mechanism. In technical terms, this means that there exist pairs of neighboring datasets for which the privacy loss random variable of our mechanism is distributed as $\mathcal{N}(\mu^{2}/2,\mu^{2})$. That is also the case for the standard Gaussian mechanism. We refer to [Ste22, Proposition 3] for a discussion of the privacy loss distribution of additive Gaussian noise.
@@
[SU16] gave a lower bound on the sample complexity for answering $d$ counting queries under $(\varepsilon,\delta)$-differential privacy. They show that the Gaussian mechanism has asymptotical optimal $\ell_{1}$ error under reasonable assumptions on $\delta$. The input domain in their paper is $\{-1,1\}^{d}$ rather than $\{0,1\}^{d}$. They consider the replacement neighboring relation and in that setting the counting query problem is equivalent for the two domains. It is easy to translate any estimate from one domain to the other using a simple post-processing step assuming that 00 should be mapped $-1$. In contrast, the problem differs between the two domains under the add/remove neighboring relation. We cannot directly translate an estimate from one domain to the other using post-processing since the size of datasets is not fixed. However, the representation of the problem in the domain $[-1,1]$ plays a key role in the privacy proof for our mechanism.
@@
The mechanism we present in this paper adds unbiased correlated Gaussian noise designed to fit the structure of the queries. Adapting Gaussian noise to the shape of the sensitivity space is not a novel concept. We refer to mechanisms that add (possibly correlated) unbiased noise as general Gaussian mechanisms. [NTZ16] gave an algorithm for answering any set of linear queries under approximate differential privacy. Linear queries are a broad class of queries that includes all problems considered in this paper. They give an algorithm for choosing the shape of Gaussian noise for any linear queries by fitting an ellipse to the convex hull of the sensitivity space. [NT24] considered the problem of mean estimation under the replacement neighborhood when data points are from some bounded domain. They show that the general Gaussian mechanism achieves near-optimal error among all unbiased mechanisms for a broad class of error measures. However, their algorithm for computing the shape of noise for arbitrary queries can be inefficient.
@@
In contrast to the general results of [NTZ16, NT24], we only consider a subset of linear queries whose sensitivity space has a certain structure. This allows us to tailor our mechanism specifically towards that structure. Concurrent with our work, [JY24] gave efficient constructions for adding elliptical Gaussian noise to two types of linear queries. They present closed-form solutions for the shape of the optimal ellipse for the Count and Vote queries. Most relevant to this paper is the Count problem which we discuss in more detail in Section 4.1. They consider the problem where entries are in $[0,b]$ rather than $[0,1]$. Note that the problem for those settings are equivalent as we can simply scale the input by a factor of $b$.
@@
[LT23] gave an $(\varepsilon,\delta)$-differentially private mechanism for releasing a Misra-Gries sketch. At a high level, their mechanism is very similar to ours. They exploit the structure of the sensitivity space of the Misra-Gries sketch and add much less noise compared to standard techniques that scale the noise by the $\ell_{1}$ or $\ell_{2}$ sensitivity. They add the same noise sample to all counters in the sketch to handle a special case where all counters differ by $1$ between some pairs of neighboring datasets. This allows them to reduce the magnitude of noise for each counter significantly because Misra-Gries sketches for neighboring datasets only differ by $1$ in a single counter except for this special case. They present some intuition behind the privacy guarantees in the technical overview of their paper using a $d+1$ representation similar to the one we use in Lemma 8. However, the structure of the sensitivity space for the queries we consider is different than the structure for the Misra-Gries sketch.
@@ 6 Conclusion and future work
In this paper we presented a simple variant of the general Gaussian mechanism for the fundamental task of answering $d$ independent queries in $[0,1]$ under the add/remove neighboring relation. In Section 4 we extended our mechanism to a setting where queries are divided into $m$ groups that form a partition of all queries and each data point only contributes to a single group. The general idea of our work is to take advantage of the hierarchical structure of these queries. Adding the same noise sample to all queries in a group allows us to reduce the noise added independently to each query. We believe this technique can be used to design differentially private mechanisms for other problems with similar structures. For example, we could consider settings where the groups of queries overlap instead of forming a partition. Alternatively, data points might contribute to a bounded number of groups instead of only one. In such settings we might need to increase the magnitude of correlated noise. Another interesting avenue to explore is whether correlated noise can improve constants for the Gaussian Sparse Histogram Mechanism [WKZK24]. We presented our work as a variant of the Gaussian mechanism, but a similar idea can be applied to other mechanisms such as the Laplace mechanism for pure differential privacy.
@@ Acknowledgments
We thank Joel Daniel Andersson and other members of the Providentia project at University of Copenhagen for helpful discussions. We thank anonymous reviewers for suggestions that helped improve the paper.

### v3

@@ 4.1 Bounded density
Figure 3: Comparison of error for the bounded Count problem for $d=1000$. Note that the optimal Count ellipse by [JY24] requires $k\leq d/2$. The left plot corresponds to the right plot in [JY24, Figure 1] where we included our mechanism for $k>d/2$. The line for the $\ell_{2}$ Ball represents the standard Gaussian mechanism in this setting.
Here we consider a setting where each data point contributes to a bounded number of queries. This setting was recently studied by [JY24]. They gave efficient constructions for adding elliptical Gaussian noise to two classes of queries. The relevant query for this paper is the Count problem, which is equivalent to the setting from Section 3 with the additional constraint that each data point has at most $k\leq d$ non-zero entries. Under this restricted setting the standard Gaussian mechanism adds noise to each query where the variance scales only by $k$ rather than $d$. Similar to Algorithm 1, the mechanism by [JY24] adds additional noise along ${\bf 1}^{d}$. They give closed-form solutions for the shape of the optimal ellipse minimizing expected squared $\ell_{2}$ norm. However, their result only applies when $k\leq d/2$. The setting we considered in Section 3 corresponds to the case where $k=d$. A natural question is whether the error can be improved for the range $d/2<k<d$. It turns out that there is almost no room for improvement. In fact, our mechanism is almost optimal for any $k\geq d/2$ as we show below.
In Figure 3 we recreated a plot from Figure 1 of [JY24]. They plot the error relative to the standard Gaussian mechanism. The improvement increases with $k$, but their mechanism is undefined for $k>d/2$. We extended the line with our mechanism for $k>d/2$ showing the improvement from elliptical noise for all values of $k$. We also plot all three mechanisms without normalizing the variance. From that plot it is clear that their mechanism should be used for small $k$ while our mechanism is preferred for any $k>d/2$. Next, we bound the difference in variance between our mechanism and the optimal general Gaussian mechanism for any $d/2\leq k\leq d$.
We first restate a result of [JY24] since the variance of their mechanism is optimal among Gaussian mechanisms for any $k\leq d/2$. Here we are interesting in the variance of their mechanism with the least restrictive parameter, that is $k=\lfloor d/2\rfloor$. They do not explicitly state the variance in their paper, but we can calculate it from their closed form expressions describing the shape of the Count ellipse. For simplicity of presentation we only consider even values of $d$. The equation for the variance of their mechanism when $d$ is odd (and $k=(d-1)/2$) is more complicated, and we do not need it to show that our mechanism is close to optimal for any $d$.
Lemma 4.3
The Count ellipse mechanism by [JY24] with parameters $k=d/2$ (where $d$ is even) and $\rho=\mu^{2}/2$ satisfies $\mu$-GDP. The mechanism estimates each query with error distributed as
$\mathcal{N}\left(0,\frac{d+2\sqrt{d-1}}{4\mu^{2}}\right)\,.$
  • Proof.
    We do not prove the privacy guarantees of their mechanism and we only sketch the proof for the variance. For more rigorous details about the mechanism we refer to their paper and references within. For simplicity, we assume without loss of generality that $\mu=1$. In the proof of [JY24, Theorem 9.9] we find their closed form expressions for variables $a_{1}$ and $a_{2}$. We do not repeat the expressions here and refer interested readers to their paper. The mechanism adds zero-mean Gaussian noise in the shape of an ellipse with an axis along $\mathbf{1}^{d}$ of length $a_{1}$. The remaining $d-1$ axis lengths are all equal to $a_{2}$. The expected squared $\ell_{2}$ norm of the noise is $a_{1}^{2}+(d-1)\cdot a_{2}^{2}$, and the variance of the error for each query is $(a_{1}^{2}+(d-1)\cdot a_{2}^{2})/d$. By setting $k=d/2$ in the expressions for $a_{1}$ and $a_{2}$ and simplifying we have that $(a_{1}^{2}+(d-1)a_{2}^{2})/d=(d+2\sqrt{d-1})/4$. Here we left out the steps to simplify the expression for brevity.
Recall from Lemma 3.1 that the variance of Algorithm 1 is $(d+2\sqrt{d}+1)/(4\mu)$. The variance achieved by [JY24] is optimal among general Gaussian mechanisms for any $k\leq\lfloor d/2\rfloor$. The variance of an optimal mechanism is non-decreasing in $k$. As such, the variance of our mechanism is almost optimal for any $k\geq d/2$.
Lemma 4.4
Let $OPT$ denote the minimum variance for any general Gaussian mechanism under $\mu$-GDP for any $d$ and parameter $k\geq d/2$. Then the variance of Algorithm 1 is at most $OPT+2\sqrt{3}/(4\mu^{2})<OPT+0.87/\mu^{2}$.
  • Proof.
    We first consider even values of $d$ and the gap to the optimal variance for $k=d/2$. We then bound the error when $d$ is odd. If $d$ is even the difference in variance between our mechanism and the optimal Count ellipse is
    $\frac{d+2\sqrt{d}+1}{4\mu^{2}}-\frac{d+2\sqrt{d-1}}{4\mu^{2}}=\frac{1+2\sqrt{d}-2\sqrt{d-1}}{4\mu^{2}}\,.$
    This function is decreasing in $d$ and as such it is maximized for $d=2$ where $(1+2\sqrt{2}-2\sqrt{2-1})/(4\mu^{2})\leq 0.46/\mu^{2}$. Since the variance for the Count ellipse mechanism with parameter $k=d/2$ is a lower bound on $OPT$ for any $k\geq d/2$, this is an upper bound on the gap to the optimal variance for all even $d$.
    When $d>2$ is odd (in the special case of $d=1$ the variance is optimal) we can bound the error using the triangle inequality. The difference in variance of Algorithm 1 for $d$ and $d-1$ is
    $\frac{d+2\sqrt{d}+1}{4\mu^{2}}-\frac{(d-1)+2\sqrt{(d-1)}+1}{4\mu^{2}}=\frac{1+2\sqrt{d}-2\sqrt{d-1}}{4\mu^{2}}\,.$
    The value of $OPT$ when $d$ is odd and $k>d/2$ is lower bounded by the value of $OPT$ for parameters $d-1$ and $k=(d-1)/2$ that we discussed above. As such, the variance of Algorithm 1 is upper bounded by
    $OPT+\frac{1+2\sqrt{d-1}-2\sqrt{(d-1)-1}}{4\mu^{2}}+\frac{1+2\sqrt{d}-2\sqrt{d-1}}{4\mu^{2}}=OPT+\frac{2+2\sqrt{d}-2\sqrt{d-2}}{4\mu^{2}}\,.$
    This gap for the bound is again decreasing in $d$ and the bound is exact for the degenerate case of $d=3$ and $k=1$. When $k=1$ the optimal mechanism is the standard Gaussian mechanism which adds spherical Gaussian noise with variance $1/\mu^{2}$. We have that Algorithm 1 achieves variance $(4+2\sqrt{3})/(4\mu^{2})$ for $k=d=3$.
We leave finding the exact value of $OPT$ for $k>d/2$ as an open problem. As a small note, we point out that the mechanism of [JY24] does not include an estimate of the dataset size. The optimal ellipse might differ slightly if we want to return an estimate of $n$ similar to Algorithm 1. It is also worth noting that it is likely possible to slightly reduce the variance even for $k=d$ if we do not care about the estimate of $n$. Notice that there is a small gap between the corners of the sensitivity space and our ellipse in Figure 1. If we instead add Gaussian noise in the shape of a tighter ellipse we would reduce the variance slightly. We do not consider that in this paper because our mechanism is simpler, and we can return an estimate of the dataset size at almost no cost for larger $d$ as shown above in Lemma 4.4. Any improvement in variance will only be noticeable for tiny values of $d$.
@@ 5 Related work
The Gaussian mechanism satisfies several definitions of differential privacy. Exact bounds are known of $(\varepsilon,\delta)$-DP [BW18], $(\alpha,\varepsilon)$-RDP [Mir17], $\rho$-zCDP [BS16], and $\mu$-GDP [DRS22]. Our privacy guarantees follow from a reduction to the standard Gaussian mechanism. As such, our results apply to any of the above definitions of differential privacy. We choose to present our mechanism in terms of $\mu$-GDP because this definition exactly describes the privacy guarantees of the Gaussian mechanism. For all problems we consider in this paper our privacy guarantees for worst-case pairs of neighboring datasets exactly match those of the standard Gaussian mechanism. In technical terms, this means that there exist pairs of neighboring datasets for which the privacy loss random variable of our mechanism is distributed as $\mathcal{N}(\mu^{2}/2,\mu^{2})$. That is also the case for the standard Gaussian mechanism. We refer to [Ste22, Proposition 3] for a discussion of the privacy loss distribution of additive Gaussian noise.
@@
[SU16] gave a lower bound on the sample complexity for answering $d$ counting queries under $(\varepsilon,\delta)$-differential privacy. They show that the Gaussian mechanism has asymptotical optimal $\ell_{1}$ error under reasonable assumptions on $\delta$. The input domain in their paper is $\{-1,1\}^{d}$ rather than $\{0,1\}^{d}$. They only consider the replacement neighboring relation and in that setting the counting query problem is equivalent for the two domains. It is easy to translate any estimate from one domain to the other using a simple post-processing step assuming that 00 should be mapped $-1$. In contrast, the problem differs between the two domains under the add/remove neighboring relation. We cannot directly translate an estimate from one domain to the other using post-processing since the size of datasets is not fixed. However, the representation of the problem in the domain $[-1,1]$ plays a key role in the privacy proof for our mechanism.
@@
The mechanism we present in this paper adds unbiased correlated Gaussian noise designed to fit the structure of the queries. Adapting Gaussian noise to the shape of the sensitivity space is not a novel concept. We refer to mechanisms that add (possibly correlated) unbiased noise as general Gaussian mechanisms. [NTZ16] gave an algorithm for answering any set of linear queries under approximate differential privacy. Linear queries are a broad class of queries that includes all problems considered in this paper. They give an algorithm for choosing the shape of Gaussian noise for any linear query by fitting an ellipse to the convex hull of the sensitivity space. [NT24] considered the problem of mean estimation under the replacement neighborhood when data points are from some bounded domain. They show that the general Gaussian mechanism achieves near-optimal error among all unbiased mechanisms for a broad class of error measures in that setting. However, their algorithm for computing the shape of noise for arbitrary queries can be inefficient.
@@
In contrast to the general results of [NTZ16, NT24], we only consider a subset of linear queries whose sensitivity space has a certain structure. This allows us to tailor our mechanism specifically towards that structure. Concurrent with our work, [JY24] gave efficient constructions for adding elliptical Gaussian noise to two types of linear queries. They present closed-form solutions for the shape of the optimal ellipse for Count and Vote queries. Most relevant to this paper is the Count problem which we discuss in more detail in Section 4.1. They consider the problem where entries are in $[0,b]$ rather than $[0,1]$. Note that the task of privately estimating the sum is equivalent for both settings as we can simply scale the input by a factor of $b$.
@@
Concurrent to this work, [FHSS24] also considered the task of privately estimating sums and dataset size, with the goal of outputting a private mean. Their technique also relies on an embedding of the data points using an additional dimension. Similar to our mechanism, they return a private estimate of the dataset size at no additional privacy cost. They explore multiple settings in their paper. In the case of Gaussian noise for one dimensional data ($x_{i}\in[0,1]$), the error of their mechanism is identically distributed to Algorithm 1. Although highly subjective, their technique is arguably preferred when $d=1$ because their privacy proof is simple and very intuitive (see Figure 1 of their paper). However, they do not reduce the error for the sums when $d>1$, which is the main contribution of our work.
@@
[LT23] gave an $(\varepsilon,\delta)$-differentially private mechanism for releasing a Misra-Gries sketch. At a high level, their mechanism is very similar to ours. They exploit the structure of the sensitivity space of the Misra-Gries sketch and add much less noise compared to standard techniques that scale the noise by the $\ell_{1}$ or $\ell_{2}$ sensitivity. They add the same noise sample to all counters in the sketch to handle a special case where all counters differ by $1$ between some pairs of neighboring datasets. This allows them to reduce the magnitude of noise significantly because Misra-Gries sketches for neighboring datasets only differ by $1$ in a single counter except for this special case. They present some intuition behind the privacy guarantees in the technical overview of their paper using a $d+1$ representation which is similar to the one we use in Lemma 3.3. However, the structure of the sensitivity space for the queries we consider is different than the structure for the Misra-Gries sketch.
@@ 6 Conclusion and future work
In this paper we presented a simple variant of the general Gaussian mechanism for the fundamental task of answering $d$ independent queries in $[0,1]$ under the add/remove neighboring relation. In Section 4 we extended our mechanism to a setting where queries are divided into $m$ groups that form a partition of all queries and each data point only contributes to a single group. The general idea in this work is to take advantage of the hierarchical structure of these queries. Adding the same noise sample to all queries in a group allows us to reduce the noise added independently to each query. We believe this technique can be used to design differentially private mechanisms for other problems with similar structures. For example, we could consider settings where the groups of queries partially overlap instead of forming a partition. Alternatively, data points might contribute to a bounded number of groups instead of only one. In such settings we might need to increase the magnitude of correlated noise. Another interesting avenue to explore is whether correlated noise can improve constants for the Gaussian Sparse Histogram Mechanism [WKZK24]. Finally, it is worth noting that we presented our work as a variant of the Gaussian mechanism, but the technique can be adapted to other mechanisms such as the Laplace mechanism for pure differential privacy.
@@ Acknowledgments
This work was carried out when the author was at the IT University of Copenhagen. We thank Joel Daniel Andersson and other members of the Providentia project at University of Copenhagen for helpful discussions. We thank anonymous reviewers for suggestions that helped improve the paper.